# R+X: Retrieval and Execution from Everyday Human Videos

### v1

@@ Abstract
We present R+X, a framework which enables robots to learn skills from long, unlabelled, first-person videos of humans performing everyday tasks. Given a language command from a human, R+X first retrieves short video clips containing relevant behaviour, and then executes the skill by conditioning an in-context imitation learning method on this behaviour. By leveraging a Vision Language Model (VLM) for retrieval, R+X does not require any manual annotation of the videos, and by leveraging in-context learning for execution, robots can perform commanded skills immediately, without requiring a period of training on the retrieved videos. Experiments studying a range of everyday household tasks show that R+X succeeds at translating unlabelled human videos into robust robot skills, and that R+X outperforms several recent alternative methods. Videos are available at https://www.robot-learning.uk/r-plus-x.
**footnotetext: Equal contribution, random order.**footnotetext: Contact at: {g.papagiannis21, n.di-palo20}@imperial.ac.uk
@@
Figure 1: Given a language prompt, R+X first retrieves short relevant video clips extracted from a long unlabelled video of a human performing everyday tasks, recorded with a wearable camera. With the use of the retrieved video clips and a VLM, R+X performs in-context imitation learning allowing it to immediately generate and execute the desired behaviour on the robot.
1 Introduction
Robot learning of diverse, everyday tasks in natural environments, is a significant challenge. Imitation learning is a promising solution which has been widely adopted in recent years [1, 2, 3], but it remains difficult to scale up due to the cost of hardware and human demonstration time. If robots could instead learn from data which does not require any robot hardware or dedicated human time, we may be able to scale up data collection and learning substantially.
@@ 1 Introduction
In this work, we study the problem of learning robot skills from long, unlabelled, first-person videos of humans performing everyday tasks in everyday environments. "Long, unlabelled videos" means that a human simply goes about their everyday life and passively records videos of themselves, without the need to specify which behaviour is being performed. "Everyday environments" means that videos contain diversity in tasks, scenarios, objects, distractors, illumination, and camera viewpoints and motion. It is likely that such videos will become abundant through the adoption of wearable devices such as AR headsets and glasses [4, 5, 6, 7], and thus this offers a significant opportunity for future scalability if robots could learn from such videos.
@@
Previous approaches to learning from videos of humans have often relied on a set of strong constraints and assumptions, such as human videos manually aligned with robot videos, human videos manually labelled with language descriptions, or demonstrations on robot or mocap hardware [8, 9, 1, 10, 11, 12, 13]. In this work, we remove all of these constraints, and present a framework that requires only an unlabelled first-person video of a human performing tasks.
@@
Our framework, R+X, is a two-stage pipeline of Retrieval and EXecution (see Figure 1), which uses Foundation Models for both stages. Upon receiving a language command from a human, a Vision Language Model retrieves all clips where the human executes the specified task. By extracting the trajectories of the human’s hand in each clip, we then employ a few-shot in-context imitation learning method to condition on these trajectories, which enables a robot to ingest, learn from, and replicate the retrieved behaviours, whilst adapting to previously unseen settings and objects. Recent literature has demonstrated that, by finetuning large VLM on robotics data, they can transfer their common knowledge and ability to reason and plan to robotics settings [14, 15]. This, however, requires very expensive finetuning of often intractably large models. With our proposed framework, we can equally leverage these abilities, but now via retrieval and video understanding, thus without the need for any finetuning. Rather than requiring explicit policy training, the retrieved skills are executed immediately following the language command. We later demonstrate the benefits that this brings over spatial and language generalisation, compared to large monolithic policy networks.
@@
In summary, the two main properties of R+X are: 1) it enables robots to execute everyday tasks from language commands, given long, unlabelled, first-person videos, collected naturally by recording a human’s daily activities; and 2) it achieves this without the need for any training or finetuning of models, allowing a robot to learn and execute tasks immediately after the language command.
2 Related Work
@@
Collecting Robotics Data. Scaling data collection has proven to be a successful path towards increasingly more general machine learning models [16, 17, 18, 19]. To collect robotics data, a popular paradigm is through kinesthetic teaching or teleoperation, collecting datasets of paired observations and actions [3]. This, however, needs dedicated robot hardware and optionally teleoperation hardware, and needs human operators to allocate their time to proactively teach a robot new tasks. In our framework, humans interact with their environments much more naturally, collecting data passively during their daily activity, resulting in a more scalable and time efficient paradigm.
@@
Learning from Human Videos. Many recent works have proposed solutions to teach new skills to robots by observing a human executing such tasks, known as Learning from Observation. However, as we illustrate in Fig. 2, they often rely on a set of assumptions and constraints that limit their use "in-the-wild". Bahl et al. [9], Wang et al. [20] require a combination of human data and either robot exploration or teleoperation, therefore needing active assistance by humans in teaching the robot, and videos are recorded from a fixed, third-person camera. R+X relies entirely on the videos recorded by the human with a mobile camera in their natural environments. Heppert et al. [10], Zhu et al. [12], Bahety et al. [11] can learn robot skills entirely from human videos. Heppert et al. [10], Zhu et al. [12], however, they focus on replicating the object trajectory from the demonstration, and therefore cannot perform non prehensile tasks that do not involve grasping and moving, such as pushing or pressing, which R+X can execute. Bahety et al. [11] can learn a large repertoire of skills, but still relies on learning single tasks in isolation from a fixed camera, while R+X can replicate tasks from a given language command after receiving a single, long, unlabelled video depicting tens of tasks, without the need to specify which clip demonstrated which behaviour. Methods like Wang et al. [13] allow a human to naturally interact with their environment while collecting data that can teach a robot new skills. However, they require additional hardware to wear, like specialised MoCap gloves. R+X only needs a single RGB-D camera, and does not require any extrinsics calibration. This means that data could be recorded using a wearable camera, smart glasses, AR visors, and more.
@@
Language-Conditioned Policy Learning. Robots should learn a wide repertoire of skills, and be able to execute many different tasks when deployed. Language is considered a viable way to instruct robots and guide their task execution [15, 14, 21]. The common approach is to train a large, language-conditioned and image-conditioned policy that can, at test time, receive human commands and camera observations [3]. This, however, presents some problems: training such models can require enormous computational effort [14], and yet robots should ideally continually learn new skills over time, where re-training or finetuning such networks becomes problematic. R+X does not train or finetune any networks, but takes advantage of pre-trained models able to retrieve examples of relevant tasks [17], enabling robots to learn and execute new behaviour directly at deployment via in-context learning [22].
@@
3 R+X: Retrieval and Execution
Figure 4: The visual 3D keypoints of the first frame of each of the Z videos obtained from the retrieval phase along with each extracted hand joint trajectory are used as context for KAT [22], a recent in-context imitation learning method. To execute a skill, visual 3D keypoints are extracted from the live observation and used as input to an imitation learning model which generates a sequence of hand joints. By mapping this sequence to gripper poses the robot executes the desired task.
@@ 3 R+X: Retrieval and Execution
We now describe R+X, our proposed method to learn robot skills from long, unlabelled videos of humans interacting with their environments. We assume R+X has access to a long, unlabelled video of a human performing a multitude of tasks in many different locations. We call this long, first-person video of everyday tasks the "human video" $\mathcal{H}$ in the rest of the paper. The goal of our method is to learn to emulate the behaviours recorded in the human video upon receiving a language command from the human.
@@
Overall, from a high-level perspective, our pipeline takes three inputs during deployment: 1) a single, long, unlabelled video of all the interactions recorded by the humans, $\mathcal{H}$; 2) a task to execute in language form, $\mathcal{L}$; and 3) the robot’s current observation as an RGB-D image, $\mathcal{O}_{live}$. It then outputs a trajectory of 6-DoF gripper poses, which are executed by the robot to perform the commanded task.
@@
There are however a set of non-trivial challenges. While the robot receives a language command at deployment, it receives no language information before deployment: it only has access to the unlabelled video. Additionally, the robot receives no action information: unlike the case of teleoperation, no joints or end-effector position/velocity/acceleration data are recorded, and the correct movements need to be inferred by the videos alone; this problem is additionally complex due to the cross-embodiment between human videos and robot actions. Furthermore, as the human is interacting with their natural environment, the visual observations can be filled with distractor objects typical of everyday environments, which are unrelated to the task the human is performing. To tackle all these aforementioned challenges, we leverage the abilities of Foundation Models.
@@
Specifically, at the first phase of R+X, which we call the retrieval phase, we use a VLM to extract from the human video all the examples of the desired behaviour described in the prompt $\mathcal{L}$, as a list of $Z$ shorter video clips $[\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{Z}]$. We map the $Z$ videos into a lower dimensional 3D representation, extracting for each video: (1) a list of $K$ visual 3D keypoints that describe the scene, $[k_{1},k_{2},\dots,k_{K}]$ where $k=(x_{k},y_{k},z_{k})$, and (2) the movement of the human’s hand as a trajectory of length $T$ of 21 hand joints that parametrise the MANO hand model [23], $\mathcal{J}=[j_{1},j_{2},\dots,j_{T}]$, where $j=[[x_{0},y_{0},z_{0}],\dots,[x_{21},y_{21},z_{21}]]$. Finally, at the second stage of R+X, which we call the execution phase, to emulate the behaviours observed in the retrieved video clips, we condition a few-shot in-context imitation learning model [22] on this data which, given a live observation of the environment, generates a trajectory of 3D hand joints to execute the desired behaviour described in the prompt. To map such joints to gripper poses, we designed a heuristic that we describe in the Supplementary Material.
@@
3.1 Retrieval: Extracting visual examples from a long, unlabelled video
@@
The first main phase of R+X is retrieval, shown in Figure 3. Upon receiving a language command $\mathcal{L}$, and given the human video $\mathcal{H}$, the goal is to retrieve all video clips from the human video that depict the execution of the requested task. This is accomplished using a recent VLM, Gemini Pro 1.5 Flash [24]. Gemini, which we denote $\mathcal{G}$, is natively multi-modal and can take as input images, videos, and text, and outputs text. We prompt the model with the human video, and ask it to retrieve the starting and ending seconds at which the received task happens. The inputs of this phase are therefore a language command and the human video, and the output is a list of $Z$ shorter video clips demonstrating the desired task, $\mathcal{G}(\mathcal{H},\mathcal{L})\rightarrow[\mathcal{V}_{1},\dots,\mathcal{V}_{Z}]$. Each clip comprises $T$ RGB-D frames, where $T$ can vary across different clips.
3.1.1 Preprocessing videos into a sparse 3D representation
@@
Figure 5: Visual keypoints are extracted autonomously from RGB via a matching algorithm [26].
@@
We first compute the DINO descriptors for the first frame of each of the $Z$ videos, obtaining a list of $Z$ different $N\times 384$ outputs, where $N$ is the number of patches of each image [27, 28, 25]. Via a clustering and matching algorithm [26], we select from the $N\times 384$ descriptors of the first frame of the first video, $\mathcal{O}_{\mathcal{V}_{1},1}$, a list of the $K$ descriptors that are the most common in all the remaining $Z-1$ frames, as shown in Figure 5. We denote these descriptors as $\mathcal{D}\in\mathbb{R}^{K\times 384}$. Therefore, this way, we autonomously extract descriptors that focus on the object of interest, as its appearance is common among videos, while distractors and overall scene will vary [26, 22].
@@
Finally, for each of the $K$ descriptors in $\mathcal{D}$, we extract keypoints by finding the $K$ nearest neighbours between the $N\times 384$ descriptors of each of the $Z$ frames, and compute their 2D coordinates. We then project these in 3D using each frame’s depth image and known camera intrinsics. As such, given the retrieved clips, we obtain and store a list $\Lambda=[\mathcal{K}_{V_{1}},\dots,\mathcal{K}_{V_{z}}]\in\mathbb{R}^{Z\times K\times 3}$ of visual 3D keypoints, $K$ for each clip, where each $\mathcal{K}_{V_{i}}=[k_{1},...,k_{K}]$ and $k_{j}=(x_{j},y_{j},z_{j})$. For a more detailed description please refer to Amir et al. [26].
@@
Hand Joint Actions. To extract human actions from each retrieved video clip we use the HaMeR model [29], a recent technique for 3D hand pose estimation from images based on DINO. In particular, using HaMeR we extract the 3D hand pose from each video frame $\mathcal{O}_{\mathcal{V}_{z},t}$ (where $1\leq t\leq T$) of each of the $Z$ clips, represented as a set of 21 3D points, $j_{\mathcal{V}_{z},t}$, describing the hand joints that parameterise the MANO hand model, as it is commonly done in the literature [23]. As the camera moves between frames, we design a stabilisation technique to compute transformations between camera poses that is robust to dynamic scenes (for more details please see our Supplementary Material). This enables us to express the extracted hand joints relative to a single reference frame, that of the first camera frame of each clip. For each video clip $\mathcal{V}_{z}$, this process results in a sequence of 3D hand joint actions $\mathcal{J}_{\mathcal{V}_{z}}=[j_{\mathcal{V}_{z},1},\dots,j_{\mathcal{V}_{z},T}]$, expressed relative to the first frame of each video $\mathcal{O}_{\mathcal{V}_{z},1}$ where the visual 3D keypoints are also expressed in. As such, at the end of this process we are left with a list $\mathcal{M}=[\mathcal{J}_{\mathcal{V}_{1}},...\mathcal{J}_{\mathcal{V}_{Z}}]$ of $Z$ hand joint action sequences.
@@
In summary, from the $Z$ retrieved video clips, we extract the list of hand joints actions $\mathcal{M}$, along with the list of visual 3D keypoints $\Lambda$, that will be used as context for our in-context imitation learning model, as $Z$ input-output pairs.
3.2 Execution: Few-Shot, In-Context Imitation from Video Examples

### v2

@@ Abstract
We present R+X, a framework which enables robots to learn skills from long, unlabelled, first-person videos of humans performing everyday tasks. Given a language command from a human, R+X first retrieves short video clips containing relevant behaviour, and then executes the skill by conditioning an in-context imitation learning method (KAT) on this behaviour. By leveraging a Vision Language Model (VLM) for retrieval, R+X does not require any manual annotation of the videos, and by leveraging in-context learning for execution, robots can perform commanded skills immediately, without requiring a period of training on the retrieved videos. Experiments studying a range of everyday household tasks show that R+X succeeds at translating unlabelled human videos into robust robot skills, and that R+X outperforms several recent alternative methods. Videos and code are available at https://www.robot-learning.uk/r-plus-x.
@@
Given a language prompt, R+X first retrieves short relevant video clips extracted from a long unlabelled video of a human performing everyday tasks, recorded with a wearable camera. By using the retrieved video clips and a VLM, R+X performs in-context imitation learning, allowing it to immediately generate and execute the desired behaviour on the robot.
footnotetext: Joint first authorship. The Robot Learning Lab at Imperial College London. Contact at: {g.papagiannis21, n.di-palo20}@imperial.ac.uk
I Introduction
@@
In this work, we study the problem of learning robot skills from long, unlabelled, first-person videos of humans performing everyday tasks in everyday environments. ”Long, unlabelled videos” means that a human simply goes by their everyday life and passively records videos of themselves without the need to specify which behaviour is being performed. ”Everyday environments” means that videos contain diversity in tasks, scenarios, objects, distractors, illumination, and camera viewpoints and motion. It is likely that such videos will become abundant through the adoption of wearable devices such as AR headsets and glasses [1, 2, 3, 4], and thus offer an unprecedented opportunity for future scalability if robots could learn from such videos, instead of learning exclusively from costly tele-operation.
@@
Previous approaches to learning from videos of humans have often relied on a set of strong constraints and assumptions, such as human videos manually aligned with robot videos, human videos manually labelled with language descriptions or demonstrations on robot or MoCap hardware [5, 6, 7, 8, 9, 10, 11]. In this work, we remove all of these constraints completely and present a framework that requires only an unlabelled first-person video depicting tens of tasks.
@@
Our framework, R+X, is a two-stage pipeline of Retrieval and EXecution shown in Figure R+X: Retrieval and Execution from Everyday Human Videos, which uses Foundation Models for both stages. Upon receiving a language command from a user, a Vision Language Model retrieves all clips where the human executes the specified task. By extracting the trajectories of the human’s hand in each clip, we then employ a few-shot in-context imitation learning method to condition on these trajectories, which enables a robot to ingest, learn from, and replicate the retrieved behaviours to previously unseen settings and objects. The recent literature demonstrated that, by finetuning large Vision Language Models on robotics data, they can transfer their common knowledge and ability to reason and plan to robotics settings [12, 13]. This, however, requires very expensive finetuning of often intractably large models. With our proposed framework, we can equally leverage these abilities but now via retrieval and video understanding, thus without the need for any finetuning.
@@
In summary, the two main properties of R+X are: 1) it enables robots to execute everyday tasks from language commands, given long, unlabelled, first-person videos collected naturally by recording a human’s daily activities, and 2) it achieves this without the need for any training or finetuning of models, allowing it to learn and execute tasks immediately, while still inhereting many strengths of large Vision Language Models.
II Related Work
@@
Collecting Robotics Data. Scaling data collection has been proven to be a successful path towards increasingly more general machine learning models [14, 15, 16, 17]. To collect robotics data, the most common paradigm is to teleoperate robots, collecting datasets of paired observations and actions [18]. This, however, needs dedicated teleoperation hardware, and needs human operators to allocate their time to actively teach a robot new tasks. In our framework, a human user interacts with their environments as usual, completing the tasks they wish, while a robot passively learns to emulate it, resulting in a more scalable and time efficient paradigm.
@@
Learning from Human Videos. Many recent works have proposed solutions to teach new skills to robots by observing a human executing such tasks. However, as we illustrate in Fig. 1, they often relied on a set of assumptions and constraints that limited their use ”in-the-wild”. [6, 19] require a combination of human data and either robot exploration or teleoperation, therefore needing active assistance of the user in teaching the robot, and videos are recorded from a fixed, third-person camera. R+X relies entirely on the videos recorded by the user with a mobile camera in their natural environments. [8, 10, 9] can learn robot skills entirely from human videos. [8, 10], however, focus on replicating the object trajectory from the demo, and cannot perform tasks that do not involve grasping and moving, such as pushing or pressing, that R+X can execute. [9] can learn a larger repertoire of skills, but still relies on learning single tasks in isolation from a fixed camera. Instead, R+X does not assume a fixed camera and can replicate tasks from a given language command after receiving a single, long, unlabelled video depicting tens of tasks, without the need to specify which clip demonstrated which behaviour. Methods like [11] allow a user to naturally interact with their environment while collecting data that can teach a robot new skills. However, they require additional hardware to wear, like specialised MoCap gloves. R+X only needs a single RGB-D camera, and does not require any extrinsics calibration. This means that data could be recorded using a wearable camera, smart glasses, AR visors, and more.
@@
Language-Conditioned Policy Learning. Language is considered a viable way to instruct robots and guide their task execution [18, 13, 12]. The common approach is to train a large, language and image conditioned policy that can, at test time, receive user commands and camera observations [18]. This, however, presents some criticalities: training such models can require enormous computational effort [12], and unlike other machine learning applications, robots should learn new skills over time. Therefore, re-training or finetuning such networks should be avoided. R+X therefore does not train or finetune any networks, but takes advantage of pre-trained models able to retrieve examples of tasks [15], and to learn to emulate and execute new behaviour directly at deployment via in-context learning [20].
Additional papers are discussed in the appendix.
@@
III R+X: Retrieval and Execution
@@
We now describe R+X, our proposed method to learn robot skills from long, unlabelled videos of humans interacting with their environments. Additional details accompanying our method can be found in the appendix and on our website. We assume R+X has access to a long, unlabelled video of a person performing a multitude of tasks in many different locations. We call this long, first-person video of everyday tasks the ”human $\mathcal{H}$ in the rest of the paper. The goal of our method is to learn to emulate the behaviours recorded in the human video upon receiving a language command.
@@
Overall, from a high-level perspective, our pipeline takes three inputs: 1) a single, long, unlabelled video of all the interactions recorded by the humans, $\mathcal{H}$ 2) a task to execute in language form, $\mathcal{L}$ and 3) the current observation of the robot as an RGB-D image, $\mathcal{O}_{live}$. It then outputs a trajectory of 6-DoF gripper poses, which are executed by the robot to tackle the task at hand.
@@
There are however a set of non-trivial challenges: while receiving a language command to execute at deployment, the robot receives no language information before deployment: the recorded video contains no more information than the recorded visual frames. Additionally, the robot receives no action information: unlike the case of teleoperation, no joints or end-effector position/velocity/acceleration data are recorded, and the correct movements need to be inferred by the videos alone: the problem is additionally complex due to the cross-embodiment between human videos and final robot actions. Furthermore, as the user is interacting with their natural environment, the visual observations can be filled with distractor objects typical of household and offices, unrelated to the task the user is performing. To tackle these challenges, we leverage the abilities of Foundation Models.
@@
Specifically, at the first phase of R+X, which we call the retrieval phase, we use a VLM to extract from the human video all the examples of the desired behaviour described in the prompt $\mathcal{L}$ as a list of $Z$ shorter video clips $[\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{Z}]$. We map the $Z$ videos into a lower dimensional 3D representation, extracting for each video: (1) a list of $K$ visual 3D keypoints that describe the scene, $[k_{1},k_{2},\dots,k_{K}]$ where $k=(x_{k},y_{k},z_{k})$, and (2) the movement of the user’s hand as a trajectory of length $T$ of 21 hand joints that parametrise the MANO hand model [21], $\mathcal{J}=[j_{1},j_{2},\dots,j_{T}]$, where $j=[[x_{0},y_{0},z_{0}],\dots,[x_{21},y_{21},z_{21}]]$. Finally, at the second stage of R+X, which we call the execution phase, to emulate the behaviours observed in the retrieved video clips, we condition a few-shot in-context imitation learning model on this data that, given a live observation of the environment, it generates a trajectory of 3D hand joints to execute the desired behaviour described in the prompt. To map such joints to gripper poses, we develop a custom procedure that we describe in the appendix.
Figure 3: The visual 3D keypoints of the first frame of each of the Z videos obtained from the retrieval phase along with each extracted hand joint trajectory are used as context for KAT. To execute a skill, visual 3D keypoints are extracted from the live observation and used as input to KAT which generates a sequence of hand joints. By mapping this sequence to gripper poses the robot executes the desired task.
@@
III-A Retrieval: Extracting visual examples from a long, unlabelled video
@@
The first main phase of R+X is retrieval, shown in Figure 2. Upon receiving a language command $\mathcal{L}$, and given the human video $\mathcal{H}$, the goal is to retrieve all video clips from the human video that depict the execution of the requested task. This is accomplished using a recent Vision Language Model (VLM), Gemini Pro 1.5 Flash [22]. Gemini, which we denote $\mathcal{G}$, is natively multi-modal and can take as input images, videos, and text, and outputs text. We prompt the model with the human video, and ask it to retrieve the starting and ending seconds at which the received task happens. The inputs of this phase is therefore a language command and the human video, and the output is a list of $Z$ shorter video clips demonstrating the desired task, $\mathcal{G}(\mathcal{H},\mathcal{L})\rightarrow[\mathcal{V}_{1},\dots,\mathcal{V}_{Z}]$. Each clip comprises $T$ RGB-D frames, where $T$ can vary across different clips.
III-A1 Preprocessing Videos into a Sparse 3D Representation
@@
We first compute the DINO descriptors for the first frame of each of the $Z$ videos, obtaining a list of $Z$ different $N\times 384$ outputs, where $N$ is the number of patches of each image [25, 26, 23]. Via a clustering and matching algorithm [24], we select from the $N\times 384$ descriptors of the first frame of the first video, $\mathcal{O}_{\mathcal{V}_{1},1}$, a list of the $K$ descriptors that are the most common in all the remaining $Z-1$ frames. We denote these descriptors as $\mathcal{D}\in\mathbb{R}^{K\times 384}$. Therefore, this way, we autonomously extract descriptors that focus on the object of interest, as its appearance is common among videos, while distractors and overall scene will vary [24, 20].
@@
Finally, for each of the $K$ descriptors in $\mathcal{D}$, we extract keypoints by finding the $K$ nearest neighbours between the $N\times 384$ descriptors of each the $Z$ frames, and compute their 2D coordinates. We then project these in 3D using each frame’s depth image and known camera intrinsics. As such, given the retrieved clips, we obtain and store a list $\Lambda=[\mathcal{K}_{V_{1}},\dots,\mathcal{K}_{V_{z}}]\in\mathbb{R}^{Z\times K\times 3}$ of visual 3D keypoints, $K$ for each clip, where each $\mathcal{K}_{V_{i}}=[k_{1},...,k_{K}]$ and $k_{j}=(x_{j},y_{j},z_{j})$. For a more detailed description please refer to [24].
@@
Hand Joint Actions. To extract human actions from each retrieved video clip we use the HaMeR model [27], a recent technique for 3D hand pose estimation from images based on DINO. In particular, using HaMeR we extract from each video frame $\mathcal{O}_{\mathcal{V}_{z},t}$ (where $1\leq t\leq T$) of each of the $Z$ clips the 3D hand pose, represented as a set of 21 3D points, $j_{\mathcal{V}_{z},t}$, describing the hand joints that parameterise the MANO hand model, as it is commonly done in the literature [21]. As the camera moves between frames, we design a stabilisation technique to compute transformations between camera poses that is robust to dynamic scenes (for more details please see the appendix). This enables us to express the extracted hand joints relative to a single reference frame, that of the first camera frame of each clip. For each video clip $\mathcal{V}_{z}$, this process results in a sequence of 3D hand joint actions $\mathcal{J}_{\mathcal{V}_{z}}=[j_{\mathcal{V}_{z},1},\dots,j_{\mathcal{V}_{z},T}]$, expressed relative to the first frame of each video $\mathcal{O}_{\mathcal{V}_{z},1}$ where the visual 3D keypoints are also expressed in. As such, at the end of this process we are left with a list $\mathcal{M}=[\mathcal{J}_{\mathcal{V}_{1}},...\mathcal{J}_{\mathcal{V}_{Z}}]$ of $Z$ hand joint action sequences.
@@
In summary, from the $Z$ retrieved video clips, we extracted the list of hand joints actions $\mathcal{M}$, along with the list of visual 3D keypoints $\Lambda$, that will be used as context for our in-context imitation learning model, as $Z$ input-output pairs.
III-B Execution: Few-Shot, In-Context Imitation from Video Examples